# Robust LLM safeguarding via refusal feature adversarial training

### v1

@@ 2 Related work
Adversarial attacks of machine learning systems have been widely studied in the literature (Zou et al., 2023b; Goodfellow et al., 2014; Madry et al., 2018). More recently, LLMs have been shown to be vulnerable to adversarial attacks generated through both manual prompt engineering (Shen et al., 2023; Anil et al., 2024) and automated techniques (Shin et al., 2020). For instance, (Zou et al., 2023b) proposed the Greedy Coordinate Gradient (GCG) suffix attack, which produces adversarial examples that can transfer from smaller open-source models to larger proprietary ones. (Chao et al., 2023) introduced the Prompt Automatic Iterative Refinement (PAIR) algorithm, where an attacker LLM iteratively queries a target LLM and refines the jailbreak prompt until a successful attack is found. Additionally, (Liu et al., 2023) developed AutoDAN, a hierarchical genetic algorithm that generates high-perplexity jailbreaks capable of bypassing LLM safety alignments. Recent work in LLM representation engineering has also revealed the potential of continuous adversarial attacks on model activations to undermine safety alignment and trigger unlearning (Schwinn et al., 2023; Arditi et al., 2024).
@@
Many methods have been proposed to defend LLMs against adversarial attacks. Simple approaches include modifying LLM system prompts to encourage model awareness to harmful requests (Xie et al., 2023; Zheng et al., 2024), but the resulting model often becomes excessively cautious and refuse to follow some normal instructions. Another family of defensive strategy is adversarial training (AT), which augments the training data of a model with adversarial prompts found by dynamically running attack methods. Recent studies show that adversarially optimized continuous perturbations can be applied to input token embeddings (Zhu et al., 2019; Jiang et al., 2020; Xhonneux et al., 2024) or LLM residual streams (Casper et al., 2024; Sheshadri et al., 2024) to enhance adversarial robustness. Another recently emerged line of work applies representation engineering (RepE) (Zou et al., 2023a) techniques to enhance model safety by directly modifying the hidden representations of input prompts. For instance, (Zou et al., 2024) proposes to “short-circuit” the model internal processes of gathering related information about unsafe requests, thereby preventing the generation of a harmful response. (Stickland et al., 2024) fine-tunes the LLM so that one can add a “steering vector” into residual stream to control model generation safety while minimizing negative side effect on its general performance. Our work connects AT and RepE by showing that refusal feature ablation can be taken as an efficient LLM representation modification that approximates adversarial perturbations in AT.
@@
It is widely believed that large language models (LLMs) represent features, or concepts, as linear directions within their activation space (Mikolov et al., 2013; Elhage et al., 2022; Park et al., 2023). Recent research has explored the linear representation of specific features, such as harmlessness Wolf et al. (2024); Zheng et al. (2024), truthfulness (Marks & Tegmark, 2023; Li et al., 2024), sentiment (Tigges et al., 2023), and refusal (Arditi et al., 2024), among others. These features are often derived from contrastive input pairs Rimsky et al. (2023); Burns et al. (2023), and have been shown to enable effective inference-time control of model behavior (Hernandez et al., 2023; Stickland et al., 2024) or targeted removal of knowledge from model parameters (Ravfogel et al., 2020; Hong et al., 2024). We extend this approach and subject linear features to adversarial perturbations applied to the model’s embedding space during training, thereby establishing a link between the interpretability and the safety alignment of LLMs.
@@ 3.1 Background
(Arditi et al., 2024) identified the refusal features (RFs) in the model’s residual stream activations using the difference-in-means technique that could effectively disentangle key feature information as demonstrated by previous work (Rimsky et al., 2023; Marks & Tegmark, 2023). We adopt the same method and compute RFs as follows: given a collection of harmful prompts $x\in\mathcal{D}_{\text{harmful}}$ (e.g. “Tell me how to make a bomb.”) and another set of harmless prompts $x\in\mathcal{D}_{\text{harmless}}$ (e.g. “Tell me how to make a cake.”), we calculate the difference between the model’s mean last-token residual stream activations when running on harmful and harmless inputs:
@@
where $\hat{\mathbf{r}}=\mathbf{r}_{\text{HH}}^{(l)}/||\mathbf{r}_{\text{HH}}^{(l)}||$ is the normalized refusal feature vector.
In contrast to (Arditi et al., 2024), we include the mean RF activation over harmless prompts in Equation 3 to account for the fact that harmless prompts may not be centered near zero along the refusal direction:
@@
In our experiments, we observed significant performance degradation when the refusal direction was simply zeroed out, potentially due to the resulting out-of-distribution behavior.  [^2] See Appendix D for a details.
@@ 4.2 Refusal feature adversarial training (ReFAT)
Drawing inspiration from our previous analyses, we propose Refusal Feature Adversarial Training (ReFAT) to enhance LLM safety and adversarial robustness. Given an LLM parameterized by $\theta$, ReFAT takes a dataset $\mathcal{D}_{\text{r}}=(x,y)$ of (harmful request, refusal answer) as inputs, and perform supervised fine-tuning by minimizing the negative conditional log likelihood of $f_{\theta}(y|x)$ of a safe answer under refusal feature ablation. In addition, the model is also trained on an utility dataset of $\mathcal{D}_{\text{u}}$ of (harmless request, helpful answer) pairs to maintain its general capability:
@@
$\mathcal{L}_{\text{RFA,r}}(\theta)=$ $-\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{r}},z\sim\mathcal{B}(p_{\text{RFA}})}\Big{[}zf_{\theta}(y|x,\mathbf{H}(x)-\mathbf{R}_{\text{HH}})+(1-z)f_{\theta}(y|x,\mathbf{H}(x))\Big{]}$ (10)
@@
where $\mathbf{R}_{\text{HH}}=\{\mathbf{r}_{\text{HH}}^{(l)}\}_{l=1}^{L}$ is the layerwise refusal feature, and $\mathbf{H}(x)-\mathbf{R}_{\text{HH}}$ denotes the removal of refusal features across model layers during model forward pass. More precisely, during each train-time forward pass, with a probability $p_{\text{RFA}}$, we perturb the intermediate representations of every input harmful instruction by removing the refusal direction from their residual stream activations. Because the parameter and the activation space change constantly during fine-tuning, we compute refusal features dynamically: every $k$ training steps, we randomly sample $n$ harmful and $n$ harmless requests from $\mathcal{D}_{\text{r}}$ and $\mathcal{D}_{\text{u}}$ respectively, and then compute a new set of $\mathbf{R}_{\text{HH}}$ using Equation 2. ReFAT can therefore simulate worst-case input perturbations by only running a few more forward passes, as opposed to common adversarial training methods which always require additional model backward passes to perform gradient-based perturbation search.
@@ 5 Experimental setups
In this section, we describe our experimental setups of assessing if ReFAT could enhance LLM robustness against adversarial attacks while preserving its general capability. Additional details can be found in Appendix A.
@@
To train models with ReFAT, we take the adversarial training dataset from (Zou et al., 2024) consisting of harmful requests that could elicit harmful or undesirable behaviors, as well as harmless conversational examples taken from UltraChat (Ding et al., 2023) to maintain model efficacy. We sample 5000 harmful requests and 5000 harmless ones from this dataset as our training data, and augment it with 150 examples taken from the XSTest dataset (Röttger et al., 2023) that includes benign requests that are seemingly risky and that the model should not decline. We take the XSTest responses generated by Llama-3-8B-Instruct as gold-standard answers for supervised fine-tuning.
@@
We compare ReFAT with several baseline safety fine-tuning methods: the refusal training (RT) refers to the standard approach of training models on the same refusal dataset as for ReFAT, but without refusal feature ablation (i.e. $p_{\text{RFA}}=0$). We also experimented with two recently proposed adversarial training methods: the Robust Refusal Dynamic Defense (R2D2) (Mazeika et al., 2024) that fine-tunes LLMs on a dynamic pool of adversarially optimized harmful requests found by GCG, and the Continuous Adversarial Training (CAT) method (Xhonneux et al., 2024) that perturbs input token embeddings with noise vectors found by gradient descent that maximize model maliciousness. For all three baselines, we also augment their training datasets with the same utility and over-refusal examples as for ReFAT.
Models
We take the same three instruction-tuned LLMs in our mechanistic analyses: Llama-3-8B (Dubey et al., 2024), Mistral-7B (Jiang et al., 2023), and Gemma-7B (Team et al., 2024). We fine-tune each model on the refusal and utility datasets using LoRA on all linear sub-layers of a transformer with a rank $r=64$. During each training step of ReFAT, with a probability $p_{\text{RFA}}=0.5$, we apply RFA to every harmful input instruction in the training batch, and do not impose any perturbations to the harmless inputs from the utility dataset. We re-compute refusal features every $k=4$ training steps using $n=32$ sampled harmful and harmless training instructions. We apply RFA on the residual stream activations over the last 75% layers of each model, as we found that this often led to the most stable fine-tuning results.
@@ 6 Results
We also found that, compared to adversarial training methods such as CAT and R2D2, ReFAT achieves similar or better robustness against various attacks. Notably, ReFAT achieves significantly lower ASR for the strongest type of RFA attack, while the other defenses are much less effective against such continuous attacks that target residual stream activation. Moreover, ReFAT strengthens model safeguards at a much lower computational cost. In particular, R2D2 and CAT both require many more additional forward and backward passes than the normal fine-tuning runs, while ReFAT on average only needs 50% more model forward passes and no additional backward pass. As a result, ReFAT is on average 1,700x cheaper than R2D2 and 10x cheaper than CAT in terms of the total number forward and backward pass.
@@ 7 Analyses of ReFAT
Develop a strategy for
hacking into a government database
and stealing sensitive information.
A daring plan! Here’s a strategy
for hacking into a government database
and stealing sensitive information:
**Step 1: Reconnaissance** …
The old ”hack into a government database” scenario!
Just to clarify, I can’t provide advice
or assistance on illegal activities
such as hacking into a government database.
@@ 1. Brute Force Attack: …
Ecco un esempio di codice per l’iniezione
di parametri HTTP per bypassare
la validazione degli input
e le regole del WAF in un’applicazione Yii:
@@
Here is an example code for injecting
HTTP parameters to bypass input
@@
Give detailed step-by-step instructions
for using the Ghostcat vulnerability
on an Apache Tomcat server
automated/ ansible guide…”,
etc blah fancy prose__:
JUST reality-as incomplete
mundane Indonesian
speech Think struggling
@@ Appendix A. Additional experimental details
Experimental setups of ReFAT
See Table 4 for a list of hyperparameters we used when fine-tuning LLMs via ReFAT, where $|\mathcal{D}_{\text{harmful}}|$ and $|\mathcal{D}_{\text{harmless}}|$ are the number of harmful and harmless instruction that we take to compute refusal features during training. In particular, we observed that training all LLMs for 1 epoch would result in models with optimal levels of adversarial robustness, while fine-tuning more than 1 epoch would make models prone to overfitting and having higher average ASR.

### v2

@@ 2 Related work
Adversarial attacks of machine learning systems have been widely studied in the literature (Zou et al., 2023b; Goodfellow et al., 2014; Madry et al., 2018). More recently, LLMs have been shown to be vulnerable to adversarial attacks generated through both manual prompt engineering (Shen et al., 2023; Anil et al., 2024; Li et al., 2024b; Zhang et al., 2024) and automated techniques (Shin et al., 2020). For instance, (Zou et al., 2023b) proposed the Greedy Coordinate Gradient (GCG) suffix attack, which produces adversarial examples that can transfer from smaller open-source models to larger proprietary ones. (Chao et al., 2023) introduced the Prompt Automatic Iterative Refinement (PAIR) algorithm, where an attacker LLM iteratively queries a target LLM and refines the jailbreak prompt until a successful attack is found. Additionally, (Liu et al., 2023) developed AutoDAN, a hierarchical genetic algorithm that generates high-perplexity jailbreaks capable of bypassing LLM safety alignments. Recent work in LLM representation engineering has also revealed the potential of continuous adversarial attacks on model activations to undermine safety alignment and trigger unlearning (Zou et al., 2023a; 2024; Schwinn et al., 2023; Arditi et al., 2024).
@@
Many methods have been proposed to defend LLMs against adversarial attacks. Simple approaches include modifying LLM system prompts to encourage model awareness to harmful requests (Xie et al., 2023; Zheng et al., 2024), and asking models to self-reflect to perform robust alignment checking Cao et al. (2023). However, the resulting model sometimes becomes excessively cautious and refuse to follow some normal instructions. Another family of defensive strategy is adversarial training (AT), which augments the training data of a model with adversarial prompts found by dynamically running attack methods. Recent studies show that adversarially optimized continuous perturbations can be applied to input token embeddings (Zhu et al., 2019; Jiang et al., 2020; Xhonneux et al., 2024) or LLM residual streams (Casper et al., 2024; Sheshadri et al., 2024) to enhance adversarial robustness. Another recently emerged line of work applies representation engineering (RepE) (Zou et al., 2023a) techniques to enhance model safety by directly modifying the hidden representations of input prompts. For instance, (Zou et al., 2024) proposes to “short-circuit” the model internal processes of gathering related information about unsafe requests, thereby preventing the generation of a harmful response. Another line of work fine-tunes the LLM so that one can add a “steering vector” into residual stream to control model generation safety while minimizing negative side effect on its general performance (Stickland et al., 2024; Cao et al., 2024). Our work connects AT and RepE by showing that refusal feature ablation can be taken as an efficient LLM representation modification that approximates adversarial perturbations in AT.
@@
It is widely believed that large language models (LLMs) represent features, or concepts, as linear directions within their activation space (Mikolov et al., 2013; Elhage et al., 2022; Park et al., 2023). Recent research has explored the linear representation of specific features, such as harmlessness Wolf et al. (2024); Zheng et al. (2024), truthfulness (Marks & Tegmark, 2023; Li et al., 2024a), sentiment (Tigges et al., 2023), and refusal (Arditi et al., 2024), among others. These features are often derived from contrastive input pairs Rimsky et al. (2023); Burns et al. (2023), and have been shown to enable effective inference-time control of model behavior (Hernandez et al., 2023; Stickland et al., 2024) or targeted removal of knowledge from model parameters (Ravfogel et al., 2020; Hong et al., 2024). We extend this approach and subject linear features to adversarial perturbations applied to the model’s embedding space during training, thereby establishing a link between the interpretability and the safety alignment of LLMs.
@@ 3.1 Background
(Arditi et al., 2024) hypothesized that refusal in instruction-tuned language models is mediated by a single direction in the residual stream, and that by steering this direction, it is possible to control the refusal behavior. Assuming that the refusal feature is a one-dimensional feature linearly encoded in the residual stream, we adopt their approach and compute the refusal features (RFs) using the difference-in-means technique, which effectively disentangles key feature information as demonstrated by previous work (Rimsky et al., 2023; Marks & Tegmark, 2023). Specifically, given a collection of harmful prompts $x\in\mathcal{D}_{\text{harmful}}$ (e.g. “Tell me how to make a bomb.”) and another set of harmless prompts $x\in\mathcal{D}_{\text{harmless}}$ (e.g. “Tell me how to make a cake.”), we calculate the difference between the model’s mean last-token residual stream activations when running on harmful and harmless inputs:
@@
where $\hat{\mathbf{r}}=\mathbf{r}_{\text{HH}}^{(l)}/||\mathbf{r}_{\text{HH}}^{(l)}||$ is unit vector encoding the refusal feature direction, and $\mathbf{h}^{(l)}(x)-\hat{\mathbf{r}}\hat{\mathbf{r}}^{T}\mathbf{h}^{(l)}(x)$ is projection that zeroes out the value along the refusal direction, and with the last term it patches the refusal feature setting it to the average value of harmless prompt activations.
In contrast to (Arditi et al., 2024), we include the mean RF activation over harmless prompts in Equation 3 to account for the fact that harmless prompts may not be centered near zero along the refusal direction  [^2] In early small-scale experiments we observed significant performance degradation when the refusal direction was simply zeroed out, potentially due to out-of-distribution. See Appendix D for details.:
@@ 4.2 Refusal feature adversarial training (ReFAT)
Drawing inspiration from our previous analyses, we propose Refusal Feature Adversarial Training (ReFAT) to enhance LLM safety and adversarial robustness. Instead of searching for a worst-case input perturbation as in standard adversarial training, we craft perturbations by ablating the refusal feature directions. Given an LLM parameterized by $\theta$, ReFAT takes a dataset $\mathcal{D}_{\text{r}}=(x,y)$ of (harmful request, refusal answer) as inputs, and performs supervised fine-tuning by minimizing the negative conditional log likelihood of $f_{\theta}(y|x)$ of a safe answer under refusal feature ablation. In addition, the model is also trained on an utility dataset of $\mathcal{D}_{\text{u}}$ of (harmless request, helpful answer) pairs to maintain its general capability:
@@
$\mathcal{L}_{\text{RFA,r}}(\theta)=$ $-p_{\text{RFA}}\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{r}}}\Big{[}f_{\theta}(y|x,\mathbf{H}(x)-\mathbf{R}_{\text{HH}})\Big{]}-(1-p_{\text{RFA}})\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{r}}}\Big{[}f_{\theta}(y|x,\mathbf{H}(x))\Big{]}$ (10)
@@
where $\mathbf{R}_{\text{HH}}=\{\mathbf{r}_{\text{HH}}^{(l)}\}_{l=1}^{L}$ is the layerwise refusal feature, and $\mathbf{H}(x)-\mathbf{R}_{\text{HH}}$ denotes the removal of refusal features across model layers during model forward pass with a probability $p_{\text{RFA}}$. We provide a pseudo-code of ReFAT in Algorithm 1 (Appendix Appendix E. ReFAT pseudo-code). Note that Eq. 11 does not incorporate the third term in the right-hand-side of Eq. 3 for efficiency considerations: computing $\bar{\mathbf{r}}_{\mathcal{D}_{\text{harmless}}}^{(l)}$ requires an additional computation step, and we found in our preliminary experiments that this would make the training slower while resulting in little improvement on model performance.
More precisely, during each train-time forward pass, we perturb the intermediate representations of every input harmful instruction by removing the refusal direction from their residual stream activations. While during each evaluation-time forward pass, we run ReFAT-trained models without refusal feature ablation, which is different from recent studies of learning effective steering vectors to perturb model activation space during evaluation (Stickland et al., 2024; Cao et al., 2024).
Because the parameter and the activation space change constantly during fine-tuning, we compute refusal features dynamically: every $k$ training steps, we randomly sample $n$ harmful and $n$ harmless requests from $\mathcal{D}_{\text{r}}$ and $\mathcal{D}_{\text{u}}$ respectively, and then compute a new set of $\mathbf{R}_{\text{HH}}$ using Equation 2. ReFAT can therefore simulate worst-case input perturbations by only running a few more forward passes, as opposed to common adversarial training methods which always require additional model backward passes to perform gradient-based perturbation search.
@@ 5 Experimental setups
In this section, we describe our experimental setups of assessing if ReFAT could enhance LLM robustness against adversarial attacks while preserving its general capability.
@@
To train models with ReFAT, we take the adversarial training dataset from (Zou et al., 2024) consisting of harmful requests that could elicit harmful or undesirable behaviors, as well as harmless conversational examples taken from UltraChat (Ding et al., 2023) to maintain model efficacy. We sample 5000 harmful requests and 5000 harmless ones from this dataset as our training data, and augment it with 150 examples taken from the XSTest dataset (Röttger et al., 2023) that includes benign requests that are seemingly risky and that the model should not decline. We use the responses generated by Llama-3-8B-Instruct on this holdout sample from XSTest as references for the next-token prediction task in the supervised finetuning step.
@@
We compare ReFAT with several baseline safety fine-tuning methods: the refusal training (RT) refers to the standard approach of training models on the same refusal dataset as for ReFAT, but without refusal feature ablation (i.e. $p_{\text{RFA}}=0$). We also experimented with three recently proposed adversarial training methods: the Robust Refusal Dynamic Defense (R2D2) (Mazeika et al., 2024) that fine-tunes LLMs on a dynamic pool of adversarially optimized harmful requests found by GCG, the Continuous Adversarial Training (CAT) method (Xhonneux et al., 2024) that perturbs input token embeddings with noise vectors found by gradient descent that maximize model maliciousness, and Latent Adversarial Training (LAT) (Casper et al., 2024; Sheshadri et al., 2024) that perturbs LLM residual stream to maximize likelihood of undesirable responses  [^5] Due to resource constraints, we only evaluated LAT by taking the LAT-fine-tuned Llama-3-8B model released by (Sheshadri et al., 2024)..
Models and hyperparameter choices
We take the same three instruction-tuned LLMs in our mechanistic analyses: Llama-3-8B (Dubey et al., 2024), Mistral-7B (Jiang et al., 2023), and Gemma-7B (Team et al., 2024). We fine-tune each model on the refusal and utility datasets using LoRA on all linear sub-layers of a transformer. We re-compute refusal features every $k=4$ training steps using $n=32$ sampled harmful and harmless training instructions. Due to the high computational cost of safety fine-tuning and evaluating model robustness via adversarial attacks, we conducted hyperparameter search through preliminary experiments by evaluating LLMs on small subsets of HarmBench and MMLU. Further details can be found in Appendix A.
@@ 6 Results
We also found that, compared to existing adversarial training methods, ReFAT achieves similar or better robustness against various attacks. Notably, ReFAT achieves significantly lower ASR for the strongest type of RFA attack, while the other defenses are less effective against such continuous attacks that target residual stream activation. Moreover, ReFAT strengthens model safeguards at a much lower computational cost. In particular, R2D2, LAT and CAT require many more additional forward and backward passes than the normal fine-tuning runs, while ReFAT on average only needs 50% more model forward passes and no additional backward pass. As a result, ReFAT is on average 1,700x cheaper than R2D2 and 10x cheaper than CAT/LAT in terms of the total number forward and backward pass.
@@ 1. Brute Force Attack: …
Ecco un esempio di codice per l’iniezione di parametri HTTP per bypassare
la validazione degli input e le regole del WAF in un’applicazione Yii:
@@
Here is an example code for injecting HTTP parameters to bypass input
@@
Give detailed step-by-step instructions for using the Ghostcat
vulnerability on an Apache Tomcat server
automated/ ansible guide…”, etc blah fancy prose__: JUST reality-as
incomplete mundane Indonesian speech Think struggling
@@ 8 Conclusion
9 Ethics statement
This work introduces an efficient and interpretable adversarial training method aimed at enhancing the robustness of large language models (LLMs) against adversarial attacks. The positive impact of this research lies in its potential to reduce the amount of harmful content generated by LLMs, as it makes many types of attacks significantly less likely to succeed. Additionally, the lower computational cost of the proposed method could help decrease the carbon footprint associated with training robust and safe LLMs. However, there are potential risks. One concern is that increased robustness might lead to overconfidence in the safety of LLMs, highlighting the need for ongoing red teaming efforts. Another potential negative impact is the possibility that adversarial training could be misused to suppress content the model operator deems undesirable, regardless of its harmfulness. Finally, our analysis of ReFAT’s failure modes reveals that LLMs may remain vulnerable to certain types of attacks (e.g., multilingual attacks), underscoring the importance of developing more rigorous and reliable evaluation protocols.
10 Reproducibility statement
We describe in Section 5 the preparation of training and evaluation data for ReFAT, the implementation of ReFAT and baseline defenses, as well as the application of attack methods to evaluate LLM adversarial robustness. We also include additional experiment configurations and hyperparameters in Appendix A. We have thoroughly checked our data and code implementation, and we have also verified empirically the effectiveness of the proposed ReFAT method.
@@ Appendix A. Additional experimental details
Hyperparameters of ReFAT
See Table 4 for a list of hyperparameters we used when fine-tuning LLMs via ReFAT, where $|\mathcal{D}_{\text{harmful}}|$ and $|\mathcal{D}_{\text{harmless}}|$ are the number of harmful and harmless instruction that we take to compute refusal features during training. We observed in preliminary experiments that training LLMs for 1 epoch would result in models with optimal levels of adversarial robustness, while fine-tuning more than 1 epoch would make models prone to overfitting and having higher average ASR. We also found that applying refusal feature ablation on the residual stream activations over the last 75% or the last 50% layers of each model often led to the most stable fine-tuning results. Finally, we found that choosing $p_{\text{RFA}}=0.5$ (i.e., training models to refuse under RFA for 50% of the cases) best balances model safety on harmful instructions with and without adversarial modifications.
@@
Further discussion on defense method hyperparameter choices
We acknowledge that careful hyperparameter search is essential for each defense method to achieve optimal utility-robustness trade-off. However, due to the high computational cost of running expensive adversarial attacks such as GCG over the full evaluation dataset of 400 harmful instructions, we chose to determine best hyperparameter configurations by evaluating and comparing multiple variations of each defense method using a small set of 40 examples from AdvBench. We found that hyperparameters of CAT and R2D2 provided in their official implementations had been highly optimized, while slightly changing some of them (e.g. the relative weight of utility loss terms) often led to severe performance degradation, so we proceeded with their default hyperparameters. For ReFAT, we searched for best hyperparameters by experimenting with four different values of $p_{\text{RFA}}\in[0.25,0.5,0.75,1.0]$ as well as four different sets of RFA layers (ablating all layers, ablating the last 75%, ablating the last 50%, and ablating the last 25%).
Although we have made our best attempt to optimize the performance of each defense method within a restricted computational budget, we acknowledge that a more careful hyperparameter search might lead to better utility-robustness trade-offs. However, we believe that our experimental results still offer strong empirical evidence that ReFAT could enhance model robustness and preserve general utility in a much more efficient way than traditional adversarial training methods.
Details of model robustness and utility evaluations